# SARDP_supp

### v1

@@ 2.2. Corrected Likelihood Estimator
See Appendix A.1 for detailed verifications. Based on the explicitly expressed bias in the derivatives led by $\mathbb{X}_{2}^{*}$, a natural method to accurately estimate parameter $\theta_{0}$ is to reconstruct an estimator using the bias-corrected first- and second- order derivatives. We refer to this estimator as the corrected likelihood estimator (CLE) , which is denoted as $\hat{\theta}_{\rm CLE}$. Accordingly, an iterative algorithm can be developed. At each iteration, we conduct three steps: (1) (Calculating) compute the first and second derivatives based on the observed $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$; (2) (Debiasing) perform bias correction for the derivatives; and (3) (Updating) conduct a Newton-Raphson-type iterative formula based on the corrected derivatives.
To be more specific, let $\hat{\theta}_{\rm CLE}^{(0)}=(\hat{\rho}_{\rm CLE}^{(0)},\hat{\beta}_{\rm CLE}^{(0)\top},\hat{\sigma}^{2(0)}_{\rm CLE})^{\top}\in\mathbb{R}^{p+2}$ be the initial estimator, which could be, for example, the QMLE estimator assuming that $\mathbb{X}_{2}^{*}$ as the true $\mathbb{X}_{2}$. Let $\hat{\theta}_{\rm CLE}^{(t)}=(\hat{\rho}_{\rm CLE}^{(t)},\hat{\beta}_{\rm CLE}^{(t)\top},\hat{\sigma}^{2(t)}_{\rm CLE})^{\top}\in\mathbb{R}^{p+2}$ be the estimator obtained in the $t$-th iteration, and $\hat{S}^{(t)}$, $\hat{\Omega}^{(t)}$ be $S$, $\Omega$ with the plugged-in estimator $\hat{\theta}_{\rm CLE}^{(t)}$ obtained in the $t$-th iteration. We next consider how to update $\hat{\theta}_{\rm CLE}^{(t)}$ to $\hat{\theta}_{\rm CLE}^{(t+1)}=(\hat{\rho}_{\rm CLE}^{(t+1)},\hat{\beta}_{\rm CLE}^{(t+1)\top},\hat{\sigma}^{2(t+1)}_{\rm CLE})^{\top}\in\mathbb{R}^{p+2}$ in the $(t+1)$-th iteration. We start from the parameter $\gamma=(\rho,\beta^{\top})^{\top}\in\mathbb{R}^{p+1}$ and its update $\hat{\gamma}_{\rm CLE}^{(t+1)}=(\hat{\rho}_{\rm CLE}^{(t+1)},\hat{\beta}_{\rm CLE}^{(t+1)\top})^{\top}$. It could be calculated that, for the first-order derivative, $\Delta\textbf{S}_{\rm CL}^{(t)}=\Delta\textbf{S}^{\rho\beta}_{\rm CL}(\hat{\theta}^{(t)}_{\rm CLE})=\Big{(}\lambda^{2}\lambda_{x}^{2}\hat{\beta}_{2}^{(t)\top}\hat{\beta}_{2}^{(t)}\mbox{tr}\Big{\{}(\hat{\Omega}^{(t)})^{-1}W\hat{S}^{(t)\top}(\hat{\Omega}^{(t)})^{-1}\Big{\}},{\bf 0}_{p_{1}}^{\top},\lambda_{x}^{2}\mbox{tr}\big{\{}(\hat{\Omega}^{(t)})^{-1}\big{\}}\\ \hat{\beta}^{(t)\top}_{2}\Big{)}\in\mathbb{R}^{p+1}$. Similarly, define $\Delta\textbf{H}_{\rm CL}^{(t)}=\Delta\textbf{H}_{\rm CL}^{\rho\beta}(\hat{\theta}^{(t)}_{\rm CLE})=\Big{(}\mathbf{H}_{\rm CL}^{\rho\rho(t)},\mathbf{H}_{\rm CL}^{\rho\beta(t)\top};\mathbf{H}_{\rm CL}^{\rho\beta(t)},\mathbf{H}_{\rm CL}^{\beta\beta(t)}\Big{)}\\ \in\mathbb{R}^{(p+1)\times(p+1)}$, which could be calculated with plugged in $\hat{\theta}_{\rm CLE}^{(t)}$ correspondingly. Then, $\hat{\gamma}_{\rm CLE}^{(t+1)}$ can be obtained using a corrected Newton-Raphson-type algorithm as,
@@
where ˙$\dot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})=\Big{(}\dot{\mathcal{L}}^{*\rho}(\hat{\theta}_{\rm CLE}^{(t)}),\dot{\mathcal{L}}^{*\beta\top}(\hat{\theta}_{\rm CLE}^{(t)})\Big{)}^{\top}$ and ¨$\ddot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})=\Big{(}\ddot{\mathcal{L}}^{*\rho\rho}(\hat{\theta}_{\rm CLE}^{(t)}),\ddot{\mathcal{L}}^{*\rho\beta\top}(\hat{\theta}_{\rm CLE}^{(t)});\\ \ddot{\mathcal{L}}^{*\rho\beta}(\hat{\theta}_{\rm CLE}^{(t)}),\ddot{\mathcal{L}}^{*\beta\beta}(\hat{\theta}_{\rm CLE}^{(t)})\Big{)}$. Note that ˙$\dot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})-\Delta\textbf{S}_{\rm CL}^{(t)}$ is a bias-corrected alternative for the first-order derivative ˙$\dot{\mathcal{L}}^{\gamma}(\theta)=\Big{(}\dot{\mathcal{L}}^{*\rho}(\theta),\dot{\mathcal{L}}^{*\beta\top}(\theta)\Big{)}^{\top}$, and ¨$\ddot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})-\Delta\textbf{H}_{\rm CL}^{(t)}$ is a bias-corrected alternative for the second-order derivative ¨$\ddot{\mathcal{L}}^{\gamma}(\theta)$ by (2.6)–(2.7) in the $t$-th iteration, where ¨$\ddot{\mathcal{L}}^{\gamma}(\theta)$ is similarly defined with ˙$\dot{\mathcal{L}}^{\gamma}(\theta)$. Then, in the $(t+1)$-th iteration, the estimate of $\sigma_{0}^{2}$ can be obtained by solving (2.8), which is $(\hat{\sigma}^{2}_{\rm CLE})^{(t+1)}$.
@@
     (Calculating Step) Compute ˙$\dot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})$ and ¨$\ddot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})$ with $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$ plugged into (2.5) and (2.9).
@@
To better illustrate the algorithm, the procedure of obtaining the CLE is presented in Algorithm 1. The algorithm can then be iteratively executed till convergence. This leads to the final estimator $\hat{\theta}_{\rm CLE}^{(t)}$. We establish the numerical convergence of $\hat{\theta}_{\rm CLE}^{(t)}$ in the next section.
@@ 2.3. Corrected Least Squares Estimator
The proposed CLE offers a feasible parameter-estimation method for the PSAR model. However, the calculation of the CLE involves the determinants of $S=I_{N}-\rho W$ and $\Omega=\sigma^{2}I_{N}+\lambda^{2}SS^{\top}$. This makes the estimation process computationally expensive for large-scale networks with complexity of order $O(N^{3})$. Consequently, inspired by the least squares estimation method proposed by Huang et al. (2019) for the univariate SAR model and Zhu et al. (2020) for the multivariate SAR model, we propose a corrected least squares estimator for the PSAR model as follows.
To illustrate the idea, we start with model (2.1) based on the true response $\mathbb{Y}$ and covariates $\mathbb{X}_{2}$. Consider the conditional expectation of $Y_{i}$, given the responses of all other nodes. Define $\mathcal{F}_{(-i)}=\sigma\{(X_{i^{\prime}},Y_{i^{\prime}}):i^{\prime}\neq i\}$ to be the $\sigma$-field generated by all other nodes. Then, under the assumption of a normal distributed ${\mathbb{E}}$, we have, $E\{Y_{i}|\mathcal{F}_{(-i)}\}=\mu_{i}+\sum_{j\neq i}^{N}\alpha_{ij}(Y_{j}-\mu_{j})$, where $\mu_{i}=E(Y_{i})$ and
@@
We make some remarks about this method. First, by equation (2.14), we could see that, for node $i$, only directly connected nodes (i.e., $a_{ij}+a_{ji}\neq 0$) and certain types of indirectly connected nodes with $\sum_{k}a_{ki}a_{kj}\neq 0$ are considered for the calculation of the conditional mean. These nodes will then be involved in the computation of the objective function (2.15). Thus, as long as the network is sufficiently sparse, the computation of the objective function will be easy. Second, the construction of the objective function here is inspired by the assumption that ${\mathbb{E}}$ follows a normal distribution. However, the formal statistical analysis does not rely on the normal assumption. Specific assumptions of the error term are provided in the next section.
Notably, the above mentioned method fails in scenarios requiring privacy protection because accurate observations of $\mathbb{X}$ and $\mathbb{Y}$ are necessary. However, the key challenge here is that we can only observe the noise-added responses $\mathbb{Y}^{*}$ and covariates $\mathbb{X}_{2}^{*}$. This leads to the noise-added loss function $\mathcal{L}_{\rm LS}^{*}(\gamma)=\mathcal{L}_{\rm LS}(\gamma,\mathbb{Y}^{*},\mathbb{X}^{*})$ instead of $\mathcal{L}_{\rm LS}(\gamma)$. Similar to the analysis of the CLE, we consider a corrected least squares method, while simultaneously correcting the biases introduced by the observational errors of $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$ on both the first and second derivatives. This results in the corrected least squares estimator (CLS) $\hat{\gamma}_{\rm CLS}$.
Specifically, let ˙$\dot{\mathcal{L}}^{*}_{\rm LS}(\gamma)=\big{(}\dot{\mathcal{L}}^{*\rho}_{\rm LS}(\gamma),\dot{\mathcal{L}}^{*\beta}_{\rm LS}(\gamma)^{\top}\big{)}^{\top}\in\mathbb{R}^{p+1}$ be the first-order derivative of $\mathcal{L}_{\rm LS}(\gamma)$ with respect to $\gamma$. Here, ˙$\dot{\mathcal{L}}^{*\rho}_{\rm LS}(\gamma)$ and ˙$\dot{\mathcal{L}}^{*\beta}_{\rm LS}(\gamma)$ represent the first-order derivatives of $\mathcal{L}_{\rm LS}(\gamma)$ with respect to $\rho$ and $\beta$, respectively. Define $\dot{d}_{\rho}$ and $\ddot{d}_{\rho}$ to be the first- and second-order derivatives of $d_{\rho}$ with respect to $\rho$, respectively. For ˙$\dot{\mathcal{L}}_{\rm LS}(\gamma)$, we have,
[…]
where $\mathbb{V}=S\mathbb{Y}-\mathbb{X}\beta$. Then, we can verify that the expectation of the first order with respect to $\gamma$ is $\mathbf{0}_{p+1}$. However, with noise-added responses $\mathbb{Y}^{*}$ and covariate $\mathbb{X}_{2}^{*}$, we have $E\big{\{}\dot{\mathcal{L}}_{\rm LS}^{*}(\gamma_{0})\big{\}}\neq\mathbf{0}_{p+1}$. Define $\mathbb{W}=W^{\top}S+S^{\top}W$ and recall that $\mathbb{V}^{*}=S\mathbb{Y}^{*}-\mathbb{X}^{*}\beta$. It can be verified that,
@@
To propose a Newton-Raphson-type algorithm with bias-corrected first- and second-order derivatives for $\mathcal{L}_{\rm LS}^{*}(\gamma)$, we still need to discuss the second order derivatives. Define ¨$\ddot{\mathcal{L}}_{\rm LS}(\gamma)=\big{(}\ddot{\mathcal{L}}_{\rm LS}^{\rho\rho}(\gamma),\ddot{\mathcal{L}}_{\rm LS}^{\rho\beta}(\gamma)^{\top};\ddot{\mathcal{L}}_{\rm LS}^{\rho\beta}(\gamma),\ddot{\mathcal{L}}_{\rm LS}^{\beta\beta}(\gamma)\big{)}$ to be the second-order derivative of $\mathcal{L}_{\rm LS}(\gamma)$. Then, it could be calculated that,
¨$\ddot{\mathcal{L}}_{\rm LS}^{\rho\rho}(\gamma)$ $=$ $4\mathbb{Y}^{\top}W^{\top}\Big{(}Wd_{\rho}^{2}S^{\top}+Sd_{\rho}^{2}W^{\top}\Big{)}\mathbb{V}-8\Big{(}\mathbb{Y}^{\top}W^{\top}S+\mathbb{V}^{\top}W\Big{)}d_{\rho}\dot{d}_{\rho}S^{\top}\mathbb{V}$
$+2\mathbb{Y}^{\top}W^{\top}Sd_{\rho}^{2}S^{\top}W\mathbb{Y}+2\mathbb{V}^{\top}\Big{(}Wd_{\rho}^{2}W^{\top}+S\ddot{d}_{\rho}d_{\rho}S^{\top}+S\dot{d}_{\rho}^{2}S^{\top}\Big{)}\mathbb{V},$
¨$\ddot{\mathcal{L}}_{\rm LS}^{\beta\beta}(\gamma)$ $=$ $2\mathbb{X}^{\top}Sd_{\rho}^{2}S^{\top}\mathbb{X},$ (2.19)
¨$\ddot{\mathcal{L}}_{\rm LS}^{\rho\beta}(\gamma)$ $=$ $2\mathbb{X}^{\top}\Big{(}Wd_{\rho}^{2}S^{\top}+Sd_{\rho}^{2}W^{\top}\Big{)}\mathbb{V}+2\mathbb{X}^{\top}Sd_{\rho}^{2}S^{\top}W\mathbb{Y}-4\mathbb{X}^{\top}Sd_{\rho}\dot{d}_{\rho}S^{\top}\mathbb{V}.$
Next, we can further define the difference between the expectations of second-order derivatives for $\mathcal{L}^{*}_{\rm LS}(\gamma)$ and $\mathcal{L}_{\rm LS}(\gamma)$ as,
$\Delta\mathbf{H}_{\rm LS}(\gamma_{0})=E\big{\{}\ddot{\mathcal{L}}^{*}_{\rm LS}(\gamma_{0})\big{\}}-E\big{\{}\ddot{\mathcal{L}}_{\rm LS}(\gamma_{0})\big{\}}=\left(\begin{array}[]{cc}\mathbf{H}_{\rm LS}^{\rho\rho}&\mathbf{H}_{\rm LS}^{\rho\beta\top}\\ \mathbf{H}_{\rm LS}^{\rho\beta}&\mathbf{H}_{\rm LS}^{\beta\beta}\\ \end{array}\right).$ (2.22)
@@
The verifications details of (2.23) and the expressions for it are given in Appendix A.2.
An iterative algorithm can then be developed, which can also be described in the three steps (calculating, debiasing, and updating) discussed for the CLE. Here, we have only listed the updating equation for simplicity. See Algorithm 2 for a detailed illustration. Let $\hat{\gamma}_{\rm CLS}^{(0)}=(\hat{\rho}_{\rm CLS}^{(0)},\hat{\beta}_{\rm CLS}^{(0)\top})^{\top}\in\mathbb{R}^{p+1}$ be the initial estimator and $\hat{\gamma}_{\rm CLS}^{(t)}=(\hat{\rho}_{\rm CLS}^{(t)},\hat{\beta}_{\rm CLS}^{(t)\top})^{\top}\in\mathbb{R}^{p+1}$ be the estimator obtained after the $t$-th iteration. The estimator in the ($t+1$)-th iteration can then be updated by a Newton-Raphson-type algorithm as follows,
$\hat{\gamma}_{\rm CLS}^{(t+1)}=\hat{\gamma}_{\rm CLS}^{(t)}-\Big{\{}\ddot{\mathcal{L}}_{\rm LS}^{*}\Big{(}\hat{\gamma}_{\rm CLS}^{(t)}\Big{)}-\Delta\textbf{H}_{\rm LS}^{(t)}\Big{\}}^{-1}\Big{\{}\dot{\mathcal{L}}_{\rm LS}^{*}\Big{(}\hat{\gamma}_{\rm CLS}^{(t)}\Big{)}-\Delta\textbf{S}_{\rm LS}^{(t)}\Big{\}},$
@@
     (Calculating Step) Compute ˙$\dot{\mathcal{L}}^{*}_{\rm LS}(\hat{\gamma}_{\rm CLS}^{(t)})$ and ¨$\ddot{\mathcal{L}}^{*}_{\rm LS}(\hat{\gamma}_{\rm CLS}^{(t)})$ with $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$ plugged into (2.16) and (2.19).
@@
     (Updating Step) Obtain $\hat{\gamma}_{\rm CLS}^{(t+1)}$ as
$\hat{\gamma}_{\rm CLS}^{(t+1)}$ $=$ $\hat{\gamma}_{\rm CLS}^{(t)}-\Big{\{}\ddot{\mathcal{L}}^{*}_{\rm LS}(\hat{\gamma}_{\rm CLS}^{(t)})-\Delta\textbf{H}_{\rm LS}^{(t)}\Big{\}}^{-1}\Big{\{}\dot{\mathcal{L}}^{*}_{\rm LS}(\hat{\gamma}_{\rm CLS}^{(t)})-\Delta\textbf{S}_{\rm LS}^{(t)}\Big{\}}.$
@@
In terms of the computational advantage, the CLS is observed to avoid the need for large-scale matrix-inversion calculations throughout the entire algorithm. This includes computing the first-order derivative, second-order derivative, and correction terms in equations (2.17)–(2.23). Consequently, this approach significantly reduces the computational time required compared to the CLE. The substantial computational advantages of $\hat{\gamma}_{\rm CLS}$ are further illustrated in the next section.

### v2

@@ 2.2. Corrected Likelihood Estimator
See Appendix A.1 for detailed verifications. Based on the explicitly expressed bias in the derivatives led by $\mathbb{X}_{2}^{*}$, a natural method to estimate parameter $\theta_{0}$ is to reconstruct an estimator using the bias-corrected first- and second-order derivatives. We refer to this estimator as the corrected likelihood estimator (CLE) , which is denoted as $\hat{\theta}_{\rm CLE}$. Accordingly, an iterative algorithm can be developed. At each iteration, we conduct three steps: (1) (Calculating) compute the first- and second-order derivatives based on the observed $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$; (2) (Debiasing) perform bias correction for the derivatives; and (3) (Updating) conduct a Newton-Raphson-type iterative formula based on the corrected derivatives.
To be more specific, let $\hat{\theta}_{\rm CLE}^{(0)}=(\hat{\rho}_{\rm CLE}^{(0)},\hat{\beta}_{\rm CLE}^{(0)\top},\hat{\sigma}^{2(0)}_{\rm CLE})^{\top}\in\mathbb{R}^{p+2}$ be the initial estimator, which could be, for example, the QMLE estimator assuming that $\mathbb{X}_{2}^{*}$ as the true $\mathbb{X}_{2}$. Let $\hat{\theta}_{\rm CLE}^{(t)}=(\hat{\rho}_{\rm CLE}^{(t)},\hat{\beta}_{\rm CLE}^{(t)\top},\hat{\sigma}^{2(t)}_{\rm CLE})^{\top}\in\mathbb{R}^{p+2}$ be the estimator obtained in the $t$-th iteration, and $\hat{S}^{(t)}$, $\hat{\Omega}^{(t)}$ be $S$, $\Omega$ with the plugged-in estimator $\hat{\theta}_{\rm CLE}^{(t)}$. We next consider how to update $\hat{\theta}_{\rm CLE}^{(t)}$ to $\hat{\theta}_{\rm CLE}^{(t+1)}$. We start from the parameter $\gamma_{0}=(\rho_{0},\beta_{0}^{\top})^{\top}\in\mathbb{R}^{p+1}$ and its update $\hat{\gamma}_{\rm CLE}^{(t+1)}=(\hat{\rho}_{\rm CLE}^{(t+1)},\hat{\beta}_{\rm CLE}^{(t+1)\top})^{\top}$. It could be calculated that, for the first-order derivative, $\Delta\textbf{S}_{\rm CL}^{(t)}=\Big{(}\lambda^{2}\lambda_{x}^{2}\hat{\beta}_{2}^{(t)\top}\hat{\beta}_{2}^{(t)}\mbox{tr}\Big{\{}(\hat{\Omega}^{(t)})^{-1}W\hat{S}^{(t)\top}(\hat{\Omega}^{(t)})^{-1}\Big{\}},{\bf 0}_{p_{1}}^{\top},\lambda_{x}^{2}\mbox{tr}\big{\{}(\hat{\Omega}^{(t)})^{-1}\big{\}}\hat{\beta}^{(t)\top}_{2}\Big{)}\in\mathbb{R}^{p+1}$. Similarly, define $\Delta\textbf{H}_{\rm CL}^{(t)}=\Big{(}\mathbf{H}_{\rm CL}^{\rho\rho(t)},\mathbf{H}_{\rm CL}^{\rho\beta(t)\top};\mathbf{H}_{\rm CL}^{\rho\beta(t)},\mathbf{H}_{\rm CL}^{\beta\beta(t)}\Big{)}\in\mathbb{R}^{(p+1)\times(p+1)}$, which could be calculated with plugged in $\hat{\theta}_{\rm CLE}^{(t)}$ correspondingly. Then, $\hat{\gamma}_{\rm CLE}^{(t+1)}$ can be obtained using a corrected Newton-Raphson-type algorithm as,
@@
where ˙$\dot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})=\Big{(}\dot{\mathcal{L}}^{*\rho}(\hat{\theta}_{\rm CLE}^{(t)}),\dot{\mathcal{L}}^{*\beta\top}(\hat{\theta}_{\rm CLE}^{(t)})\Big{)}^{\top}$ and ¨$\ddot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})=\Big{(}\ddot{\mathcal{L}}^{*\rho\rho}(\hat{\theta}_{\rm CLE}^{(t)}),\ddot{\mathcal{L}}^{*\rho\beta\top}(\hat{\theta}_{\rm CLE}^{(t)});\\ \ddot{\mathcal{L}}^{*\rho\beta}(\hat{\theta}_{\rm CLE}^{(t)}),\ddot{\mathcal{L}}^{*\beta\beta}(\hat{\theta}_{\rm CLE}^{(t)})\Big{)}$. Then, in the $(t+1)$-th iteration, the estimate of $\sigma_{0}^{2}$ can be obtained by solving (2.7), which is $\hat{\sigma}_{\rm CLE}^{2(t+1)}$. See Algorithm 1 for details. This leads to the final estimator $\hat{\theta}_{\rm CLE}^{(t)}$. We establish the numerical convergence of $\hat{\theta}_{\rm CLE}^{(t)}$ in the next section.
@@
     (Calculating Step) Compute ˙$\dot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})$ and ¨$\ddot{\mathcal{L}}^{*\gamma}(\hat{\theta}_{\rm CLE}^{(t)})$ based on $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$.
@@ 2.3. Corrected Least Squares Estimator
The proposed CLE offers a feasible parameter-estimation method for the PSAR model. However, the calculation of the CLE involves the determinants of $S=I_{N}-\rho W$ and $\Omega=\sigma^{2}I_{N}+\lambda^{2}SS^{\top}$. This makes the estimation process computationally expensive for large-scale networks with complexity of order $O(N^{3})$. Consequently, inspired by the least squares estimation method by Huang et al. (2019) and Zhu et al. (2020), we propose a corrected least squares estimator for the PSAR model as follows.
To illustrate the idea, we start with model (2.1) based on the true response $\mathbb{Y}$ and covariates $\mathbb{X}_{2}$. Consider the conditional expectation of $Y_{i}$, given the responses of all other nodes. Define $\mathcal{F}_{(-i)}=\sigma\{(X_{i^{\prime}},Y_{i^{\prime}}):i^{\prime}\neq i\}$ to be the $\sigma$-field generated by all other nodes. Then, under the assumption of a normally distributed ${\mathbb{E}}$, we have, $E\{Y_{i}|\mathcal{F}_{(-i)}\}=\mu_{i}+\sum_{j\neq i}^{N}\alpha_{ij}(Y_{j}-\mu_{j})$, where $\mu_{i}=E(Y_{i})$ and
@@
We make two remarks about this method. First, from equation (2.12), we see that, for node $i$, only directly connected nodes (i.e., $a_{ij}+a_{ji}\neq 0$) and certain types of indirectly connected nodes with $\sum_{k}a_{ki}a_{kj}\neq 0$ are considered for the calculation of the conditional mean. These nodes will then further be involved in the computation of the objective function (2.13). Thus, as long as the network is sufficiently sparse, the computation of the objective function will be efficient. Second, the construction of the objective function here is inspired by the assumption that ${\mathbb{E}}$ follows a normal distribution. However, the formal statistical analysis does not rely on the normal assumption. Specific assumptions of the error term are provided in the next section.
However, the key challenge here is that we can only observe the noise-added responses $\mathbb{Y}^{*}$ and covariates $\mathbb{X}_{2}^{*}$. This leads to the noise-added loss function $\mathcal{L}_{\rm LS}^{*}(\gamma)=\mathcal{L}_{\rm LS}(\gamma;\mathbb{Y}^{*},\mathbb{X}^{*})$ instead of $\mathcal{L}_{\rm LS}(\gamma)$. Similar to the analysis of the CLE, we consider a corrected least squares method, while simultaneously correcting the biases introduced by the observational errors of $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$ on both the first- and second-order derivatives. This results in the corrected least squares estimator (CLS) $\hat{\gamma}_{\rm CLS}$.
Specifically, let ˙$\dot{\mathcal{L}}^{*}_{\rm LS}(\gamma)=\big{(}\dot{\mathcal{L}}^{*\rho}_{\rm LS}(\gamma),\dot{\mathcal{L}}^{*\beta}_{\rm LS}(\gamma)^{\top}\big{)}^{\top}\in\mathbb{R}^{p+1}$ be the first-order derivative of $\mathcal{L}_{\rm LS}(\gamma)$ with respect to $\gamma$. Here, ˙$\dot{\mathcal{L}}^{*\rho}_{\rm LS}(\gamma)$ and ˙$\dot{\mathcal{L}}^{*\beta}_{\rm LS}(\gamma)$ represent the first-order derivatives of $\mathcal{L}_{\rm LS}(\gamma)$ with respect to $\rho$ and $\beta$, respectively. Define $\dot{d}_{\rho}$ and $\ddot{d}_{\rho}$ to be the first- and second-order derivatives of $d_{\rho}$ with respect to $\rho$, respectively. Define $\mathbb{W}=W^{\top}S+S^{\top}W$ and $\mathbb{V}^{*}=S\mathbb{Y}^{*}-\mathbb{X}^{*}\beta$. It can be verified that $E\big{\{}\dot{\mathcal{L}}_{\rm LS}(\gamma_{0})\big{\}}=\mathbf{0}_{p+1}$, and
@@
To propose a Newton-Raphson-type algorithm, we still need to discuss the second-order derivatives. Define ¨$\ddot{\mathcal{L}}_{\rm LS}(\gamma)=\big{(}\ddot{\mathcal{L}}_{\rm LS}^{\rho\rho}(\gamma),\ddot{\mathcal{L}}_{\rm LS}^{\rho\beta}(\gamma)^{\top};\ddot{\mathcal{L}}_{\rm LS}^{\rho\beta}(\gamma),\ddot{\mathcal{L}}_{\rm LS}^{\beta\beta}(\gamma)\big{)}$ to be the second-order derivative of $\mathcal{L}_{\rm LS}(\gamma)$. We can further define the difference between the expectations of second-order derivatives for $\mathcal{L}^{*}_{\rm LS}(\gamma)$ and $\mathcal{L}_{\rm LS}(\gamma)$ as,
$\Delta\mathbf{H}_{\rm LS}(\gamma_{0})=E\big{\{}\ddot{\mathcal{L}}^{*}_{\rm LS}(\gamma_{0})\big{\}}-E\big{\{}\ddot{\mathcal{L}}_{\rm LS}(\gamma_{0})\big{\}}=\left(\begin{array}[]{cc}\mathbf{H}_{\rm LS}^{\rho\rho}&\mathbf{H}_{\rm LS}^{\rho\beta\top}\\ \mathbf{H}_{\rm LS}^{\rho\beta}&\mathbf{H}_{\rm LS}^{\beta\beta}\end{array}\right).$ (2.18)
@@
The verification details of (2.19) and the expressions for it are given in Appendix A.2.
An iterative algorithm (i.e., Algorithm 2) can then be developed, which can also be described in the three steps (calculating, debiasing, and updating). Here, we only list the updating equation for simplicity. Let $\hat{\gamma}_{\rm CLS}^{(0)}=(\hat{\rho}_{\rm CLS}^{(0)},\hat{\beta}_{\rm CLS}^{(0)\top})^{\top}\in\mathbb{R}^{p+1}$ be the initial estimator and $\hat{\gamma}_{\rm CLS}^{(t)}=(\hat{\rho}_{\rm CLS}^{(t)},\hat{\beta}_{\rm CLS}^{(t)\top})^{\top}\in\mathbb{R}^{p+1}$ be the estimator obtained in the $t$-th iteration. The estimator in the ($t+1$)-th iteration can then be updated as follows,
$\hat{\gamma}_{\rm CLS}^{(t+1)}=\hat{\gamma}_{\rm CLS}^{(t)}-\Big{\{}\ddot{\mathcal{L}}_{\rm LS}^{*}\Big{(}\hat{\gamma}_{\rm CLS}^{(t)}\Big{)}-\Delta\textbf{H}_{\rm LS}^{(t)}\Big{\}}^{-1}\Big{\{}\dot{\mathcal{L}}_{\rm LS}^{*}\Big{(}\hat{\gamma}_{\rm CLS}^{(t)}\Big{)}-\Delta\textbf{S}_{\rm LS}^{(t)}\Big{\}},$ (2.20)
@@
     (Calculating Step) Compute ˙$\dot{\mathcal{L}}^{*}_{\rm LS}(\hat{\gamma}_{\rm CLS}^{(t)})$ and ¨$\ddot{\mathcal{L}}^{*}_{\rm LS}(\hat{\gamma}_{\rm CLS}^{(t)})$ based on $\mathbb{X}^{*}$ and $\mathbb{Y}^{*}$.
@@
     (Updating Step) Obtain $\hat{\gamma}_{\rm CLS}^{(t+1)}$ by (2.20).
@@
In terms of the computational advantage, the CLS is observed to avoid the need for large-scale matrix-inversion calculations throughout the entire algorithm. This includes computing the first-order derivative, second-order derivative, and correction terms in equations (2.14)–(2.19). Consequently, this approach significantly reduces the computational time required compared to the CLE. The substantial computational advantages of $\hat{\gamma}_{\rm CLS}$ are further illustrated in the numerical analysis.